# Learning to Reason Across Parallel Samples for LLM Reasoning

### v1

@@ Abstract
Scaling test-time compute brings substantial performance gains for large language models (LLMs). By sampling multiple answers and heuristically aggregate their answers (e.g., either through majority voting or using verifiers to rank the answers), one can achieve consistent performance gains in math domains. In this paper, we propose a new way to leverage such multiple sample set. We train a compact LLM, called Sample Set Aggregator (SSA), that takes a concatenated sequence of multiple samples and output the final answer, optimizing it for the answer accuracy with reinforcement learning. Experiments on multiple reasoning datasets show that SSA outperforms other test-time scaling methods such as reward model-based re-ranking. Our approach also shows a promising generalization ability, across sample set sizes, base model families and scales, and tasks. By separating LLMs to generate answers and LLMs to analyze and aggregate sampled answers, our approach can work with the outputs from premier black box models easily and efficiently.
@@ 1 Introduction
Recent advances in large language models (LLMs) have significantly enhanced their ability to perform complex reasoning tasks [8, 7]. Orthogonal to approaches to improve training LLMs through better learning objectives [29, 34], architectures [10, 30] or training dataset composition [11, 45], recent work [37] explores a new dimension: test-time scaling. While having the LLM fixed, by allocating more computation at inference time (e.g., through repeated sampling and majority voting), one can improve the final task performance.
In general, test-time scaling methods fall into two paradigms. Parallel scaling generates multiple reasoning paths independently and aggregates them via strategies such as majority voting or best-of-N selection [43, 40, 23]. Sequential scaling, on the other hand, iteratively refines a single solution, often through prompting-based self-reflection or by incentivizing iterative computation [7, 28, 19]. In this paper, we introduce a novel test-time scaling approach that leverages both parallel and sequential scaling. Figure 1 illustrates our approach in comparison with prior work.
@@
We conduct extensive experiments across diverse reasoning benchmarks under controlled test-time compute budgets. Results show that SSA substantially narrows the gap between actual model performance and oracle-best accuracy (pass@K), outperforming standard parallel strategies such as reward-based reranking. Moreover, we demonstrate that a compact SSA model can match the performance of reinforcement-trained larger models used in sequential scaling, suggesting the effectiveness of SSA as a lightweight way for sequential scaling. Further analysis highlights the generalization capabilities of SSA: SSA trained on one dataset for a particular model can successfully synthesize outputs from different model families and sizes across different tasks.
@@
  • Conceptually, we propose to reason over the output distribution instead of tuning the model internals. Specifically, instead of training the base LM, SSA optimizes over its sampled outputs. The base LM that produces answers remains to be a black box; SSA is trained only with sampled answers from base LMs.
@@ 2 Related Work
Recent research has established the effectiveness of increasing compute used at inference time, known as test-time scaling [8, 7, 37, 2]. A prominent approach for test-time scaling focuses on parallel scaling, which samples multiple answers independently and aggregates them into a single answer. This aggregation can be performed through majority voting [43, 41] or more sophisticated selection mechanisms. For instance, some methods prompt language models to select from one of the multiple samples [3], while others employ dedicated verifier models to score potential solutions [6, 40, 23, 21, 42] and take a weighted majority solution.
@@
Beyond naively sampling multiple answers with fixed decoding strategy, researchers have explored advanced search strategies, such as beam search [50, 46] and Monte-Carlo tree search (MCTS) [20, 47]. These search-based methods typically rely on verifiers [46] or process reward models [6, 42] to guide the decoding process, while still evaluating different rollouts independently. In contrast to these approaches that primarily assess samples in isolation, our approach learns to compare different samples jointly to determine the final answer.
Scaling test-time compute sequentially
@@
Another line of research focuses on sequential scaling, which increases compute by iteratively updating and refining a solution. One way is to prompt or train LLMs to self-refine their proposed solutions [27, 19, 32, 4] in an iterative manner. Recent work has demonstrated the possibility of incentivizing LLMs to spend more tokens in a single completion through reinforcement learning [7, 38] or by forcing LLMs to continue their reasoning chains by appending "wait" token [28]. While sequential scaling often yields performance improvements, recent studies debate on whether it outperforms simpler parallel scaling when controlling for compute [52, 14, 15]. Our approach bridges these paradigms by operating on parallel samples but treating them as a sequence rather than as isolated instances.
Training language models for reasoning
@@
More broadly, our work aims to enhance LLM performance on reasoning tasks, which has been a central pursuit in LLM development. Substantial efforts have been devoted to training LLMs as reasoning policies (distributions more likely to contain correct answers) through supervised fine-tuning on collections of chain-of-thought [1, 31, 25], or through reinforcement learning with rewards on intermediate steps [40, 42, 18] or based solely on final answer correctness [7]. Our research is more closely aligned with work on training LLMs as better verifiers for reasoning tasks and can complement solutions from any policy models. Existing methods train LLM verifiers to assess individual solutions using human annotations [6] or noisy labels derived from answer correctness [42, 16, 22]. Our approach differs in that we train LLMs to verify sequences of solutions using reinforcement learning with a verifiable reward signal, the final answer correctness.
@@ 3.1 Problem Formulation
To enable an LLM to function as SSA $LM_{\text{SSA}}$ for synthesizing the final answer, we provide a natural language instruction directing the model to “think carefully and thoroughly evaluate the proposed answer, and identify one correct answer from the proposed candidates”. Following common practice in prior work [7], we specify a structured output format, which we use to extract the final answer. Please refer to Appendix B Figure 8 for details of the prompt.
@@ 3.2 Training SSA with Reinforcement Learning
Following prior work [7], our reward also consider the format of the output specified in the prompt. Specifically, we let the reward $R$ to be 1.0 if the answer is correct; we let the reward be 0.05 if the output follows the format and is incorrect; we let the reward to be 0.0 if the output does not follow and format.
We use the Group-Relative Policy Optimization (GRPO [36]) as our optimization algorithm, which simplified value function of PPO [35] with a normalized reward from a group. For the convention of notation, we use $\pi_{\theta}$ for $LM_{\text{SSA}}$ with parameter $\theta$. For completeness, we describe it here. It maximizes:
@@
Implementation
We use a training set combining GSM8K and MATH [6, 13]. To prepare the dataset, we use Qwen2.5-7B-Instruct model as $\text{LM}_{\text{ans}}$ to generate five answers for each questions [33]. Then we concatenate the question and each answer (ordered randomly) as input context of SSA LLM. We filter out instances where the combined answers exceed 4000 tokens, or where fewer than three valid answers are available. This results a dataset with a total size of 17.4k. We set the GRPO sample size to be 8, resulting in $8*17.4k\approx 140k$ during training.
@@ 3.3 Training SSA With Supervised Finetuning
Implementation
We prompt Qwen 2.5 7B Instruct model for 5 candidate solutions per question in the GSM8K dataset. Then, we provide concatenated candidate solutions to GPT-4.1 Nano model along with the original question and ground-truth answer (The exact prompt is in Appendix B). GPT-4.1 Nano then provides a step-by-step reasoning process to identify and generate the best final answer from these candidates, yielding 7.47k training examples. Overall, this approach achieves a 96.24% match rate with the original ground-truth answers. [^1] When we inspected the remaining cases, we find many ground-truth labels are incorrect.
@@ 4.1 Experimental Settings
Datasets.
For evaluation, we use an array of commonly used math reasoning datasets: the test split of GSM8K, MATH as the in-domain evaluation sets, and AIME 2024 [26], AMC 2023, and Olympiad [12] as the test sets. We use the extracted answers and grade them against the ground truth answers, using the library that has been used in prior work [23].
Base Models for Candidate Generation ($\text{LM}_{\text{ans}}$).
We use the Qwen-2.5-Instruct model, with sizes 7B, 14B, and 32B to generate $K$ of answers. We use a decoding temperature 0.5 to construct the training and test dataset. In the training phase, we use $k=5$ to train the SSA model. In the testing phase, we sampled up to $k=20$ and evaluated the performance with $k=\{5,10,15,20\}$.
Base Models for SSA ($\text{LM}_{\text{SSA}}$).
We use the Qwen-2.5-base model with sizes 0.5B, 1.5B, and 3B. We chose the Qwen-2.5 model family due to its popularity and wide availability for PRM verifiers, making it possible to compare against existing PRM verifiers from the same model family.
Training Implementation Details.
For the training library, we use torchtune due to its efficient VRAM management [39]. We use GRPO batch sample group size 8, batch size 1, temperature 1.0, AdamW optimizer, KL coefficient 0.01, and learning rate 1e-5. We trained all experiments one epoch. For hardware, we use 8*H100 80GB for training. For shorter context, it is also possible to train with 48GB VRAM.
@@ 4.2 Comparison Systems
Rule-based Baselines.
We report pass@k, which reports the percentage of examples where any of $k$ number of solution is correct [2]. Pass@k also serves as the oracle performance assuming we have an oracle verifier. We also report the majority vote [43], counting the most frequent answer among the answer set.
Outcome Reward Model (ORM).
We use off the shelf Llama-3.1-8B ORM model trained with RLHF-Reward Modeling [48] to re-rank multiple samples, selecting the best one scored by the reward. We will call it Llama-ORM (8B). It is trained with 273k data. For ORM evaluations, it adds an evaluation token at the end of the answer and the model will provide a score.
Process Reward Model (PRM).
One prominent way to leverage multiple parallel sample is using process reward models. We use the Qwen-7B PRM (Qwen PRM) model from Zhang et al. [53] to re-rank the candidate solutions.
We note that compared to the PRM model, our SSA is trained with significantly less resources regarding both model scale (0.5B - 3B vs 7B) and the amount of training data. Specifically, Qwen-7B PRM is trained with more than 500,000 queries with 6 to 8 answers with step labels, resulting over 3 million total training data (about 20$\times$ larger than the training data size of SSA). For a more fair comparison, we also include Shepherd PRM [42], 7B model trained with 440k total step level data from their MCTS roll out. Additionally, Qwen PRM initialized from Qwen2.5-Math-7B-Instruct [49], whereas we initialized from the Qwen 2.5 base model.
@@
Universal Self Consistency (USC) [3].
USC is a prompting-based method takes a concatenation of the multiple parallel sample answers, and instructs a LM to generate the final answer. The exact prompt can be found in Appendix B. The task setting is equivalent to our SSA method, but the LM is not trained to optimize for the final answer.
Anlayzer and Varitions.
We train three types of the SSA models, one trained with the SFT objective only (SFT), one trained with SFT objective and then with the RL objective (SFT + RL), and one that is trained with RL objective only (RL). For SFT objective, we use the dataset described in Section 3.3 to finetune the model. We use learning rates 5e-5, 2e-5, and 1e-5 for 0.5B, 1.5B, and 3B correspondingly. For RL objective, we use GRPO method described in the Section 3.2. For SFT + RL objective, we use the trained model from the SFT objective, then we continue to train the RL objective on top of it.
For each variant, we present models of three different sizes, trained from Qwen-0.5B, 1.5B, and 3B base models checkpoint respectively.
@@ 5 Main Results
Table 1: Results (accuracy %) with $k=5$ candidate answers generated by Qwen2.5-Math-7B-Instruct as $\text{LLM}_{\text{ans}}$ model. Training Data Size is computed through total queries times number of responses for each query.
Method Train Data Size Datasets Avg
@@
[…]
USC w/ Qwen 7B - 74.98 61.20 6.67 47.50 28.78 43.83
LLama ORM (8B) 273k 93.1 67.40 13.33 47.50 28.64 49.99
Qwen PRM (7B) 3M+ 92.57 69.40 13.33 57.50 32.05 52.97
Shepherd PRM (7B) 440k 90.75 64.40 13.33 35.00 27.00 46.10
[…]
SSA RL (1.5B) 140k 92.49 76.60 10.00 52.50 38.72 54.06
SSA RL (3B) 140k 93.25 76.80 13.33 57.50 39.76 56.13

### v2

@@ Abstract
Scaling test-time compute brings substantial performance gains for large language models (LLMs). By sampling multiple answers and heuristically aggregate their answers (e.g., either through majority voting or using verifiers to rank the answers), one can achieve consistent performance gains in math domains. In this paper, we propose a new way to leverage such multiple sample set. We train a compact LLM, called Sample Set Aggregator (SSA), that takes a concatenated sequence of multiple samples and output the final answer, optimizing it for the answer accuracy with reinforcement learning. Experiments on five reasoning datasets demonstrate both the efficacy and efficiency of SSA. Notably, SSA improves over naive majority voting by 8% pass@5 on MATH. Furthermore, our 3B SSA surpasses model-based re-ranking with a much larger 72B process reward model. Our analysis also shows promising generalization ability of SSA, across sample set sizes, base model families and scales, and tasks. By separating LLMs to generate answers and LLMs to analyze and aggregate sampled answers, our approach can work with the outputs from premier black box models easily and efficiently.
@@ 1 Introduction
Recent advances in large language models (LLMs) have significantly enhanced their ability to perform complex reasoning tasks (El-Kishky et al., 2024; DeepSeek-AI et al., 2025). Orthogonal to approaches to improve training LLMs through better learning objectives (Ouyang et al., 2022; Rafailov et al., 2023), architectures (Gu & Dao, 2024; Peng et al., 2023) or training dataset composition (Gunasekar et al., 2023; Wettig et al., 2025), recent work (Snell et al., 2025) explores a new dimension: test-time scaling. While having the LLM fixed, by allocating more computation at inference time (e.g., through repeated sampling and majority voting), one can improve the final task performance.
In general, test-time scaling methods fall into two paradigms. Parallel scaling generates multiple reasoning paths independently and aggregates them via strategies such as majority voting or best-of-N selection (Wang et al., 2023; Uesato et al., 2022; Lightman et al., 2024). Sequential scaling, on the other hand, iteratively refines a single solution, often through prompting-based self-reflection or by incentivizing iterative computation (DeepSeek-AI et al., 2025; Muennighoff et al., 2025; Kumar et al., 2025). In this paper, we introduce a novel test-time scaling approach that leverages both parallel and sequential scaling. Figure 1 illustrates our approach in comparison with prior work.
@@
We conduct extensive experiments across five reasoning benchmarks under controlled test-time compute budgets. Results show that SSA substantially narrows the gap between actual model performance and oracle-best accuracy (pass@K), outperforming standard parallel strategies such as reward-based reranking. Moreover, we demonstrate that a compact SSA model can match the performance of reinforcement-trained larger models used in sequential scaling, suggesting the effectiveness of SSA as a lightweight way for sequential scaling. Further analysis highlights the generalization capabilities of SSA: SSA trained on one dataset for a particular model can successfully synthesize outputs from different model families and sizes across different tasks.
@@
  • Conceptually, we propose to reason over the output distribution. Specifically, instead of training the base LM, SSA optimizes over its sampled outputs. The base LM that produces answers remains to be a black box; SSA is trained only with sampled answers from base LMs.
@@ 2 Related Work
Scaling test-time compute in parallel. Recent research has established the effectiveness of increasing compute used at inference time, known as test-time scaling (El-Kishky et al., 2024; DeepSeek-AI et al., 2025; Snell et al., 2025; Brown et al., 2025). A prominent approach for test-time scaling focuses on parallel scaling, which samples multiple answers independently and aggregates them into a single answer. This aggregation can be performed through majority voting (Wang et al., 2023; 2024a) or more sophisticated selection mechanisms. For instance, some methods prompt language models to select from one of the multiple samples (Chen et al., 2024a), while others employ dedicated verifier models to score potential solutions (Cobbe et al., 2021; Uesato et al., 2022; Lightman et al., 2024; Li et al., 2023; Wang et al., 2024b) and take a weighted majority solution.
@@
Beyond naively sampling multiple answers with fixed decoding strategy, researchers have explored advanced search strategies, such as beam search (Yao et al., 2023; Xie et al., 2023) and Monte-Carlo tree search (MCTS) (Li et al., 2025; Xie et al., 2024). These search-based methods typically rely on verifiers (Xie et al., 2023) or process reward models (Cobbe et al., 2021; Wang et al., 2024b) to guide the decoding process, while still evaluating different rollouts independently. In contrast to these approaches that primarily assess samples in isolation, our approach learns to compare different samples jointly to determine the final answer.
@@
Scaling test-time compute sequentially. Another line of research focuses on sequential scaling, which increases compute by iteratively updating and refining a solution. One way is to prompt or train LLMs to self-refine their proposed solutions (Madaan et al., 2023; Kumar et al., 2025; Qu et al., 2024; Chen et al., 2024b) in an iterative manner. Recent work has demonstrated the possibility of incentivizing LLMs to spend more tokens in a single completion through reinforcement learning (DeepSeek-AI et al., 2025; Team et al., 2025) or by forcing LLMs to continue their reasoning chains by appending "wait" token (Muennighoff et al., 2025). While sequential scaling often yields performance improvements, recent studies debate on whether it outperforms simpler parallel scaling when controlling for compute (Zeng et al., 2025b; Hochlehnert et al., 2025a; b). Our approach bridges these paradigms by operating on parallel samples but treating them as a sequence rather than as isolated instances. The growing length of CoT traces incurs substantial inference cost. Recent efforts have sought to improve efficiency by introducing length penalties (Aggarwal & Welleck, 2025; Sui et al., 2025), adaptive thinking (Fang et al., 2025; Zhang et al., 2025b; Lou et al., 2025), or applying early stopping (Zhang et al., 2025a; Yang et al., 2025b). In contrast, our approach trains a compact aggregator that achieves both strong performance and efficiency.
@@
Training language models for reasoning. More broadly, our work aims to enhance LLM performance on reasoning tasks, which has been a central pursuit in LLM development. Substantial efforts have been devoted to training LLMs as reasoning policies (distributions more likely to contain correct answers) through supervised fine-tuning on collections of chain-of-thought (Azerbayev et al., 2024; Puerto et al., 2024; Luo et al., 2023), or through reinforcement learning with rewards on intermediate steps (Uesato et al., 2022; Wang et al., 2024b; Kazemnejad et al., 2025) or based solely on final answer correctness (DeepSeek-AI et al., 2025). Our research is more closely aligned with work on training LLMs as better verifiers for reasoning tasks and can complement solutions from any policy models. Existing methods train LLM verifiers to assess individual solutions using human annotations (Cobbe et al., 2021) or noisy labels derived from answer correctness (Wang et al., 2024b; Hosseini et al., 2024; Liang et al., 2024). Our approach differs in that we train LLMs to verify sequences of solutions using reinforcement learning with a verifiable reward signal, the final answer correctness.
@@ 3.1 Problem Formulation
To enable an LLM to function as SSA $LM_{\text{SSA}}$ for synthesizing the final answer, we provide a natural language instruction directing the model to “think carefully and thoroughly evaluate the proposed answer, and identify one correct answer from the proposed candidates”. Following common practice in prior work (DeepSeek-AI et al., 2025), we specify a structured output format, which we use to extract the final answer. Please refer to Appendix B Figure 8 for details of the prompt.
@@ 3.2 Training SSA with Reinforcement Learning
Following prior work (DeepSeek-AI et al., 2025), our reward also consider the format of the output specified in the prompt. Specifically, we let the reward $R$ to be 1.0 if the answer is correct; we let the reward be 0.05 if the output follows the format and is incorrect; we let the reward to be 0.0 if the output does not follow and format.
We use the Group-Relative Policy Optimization (GRPO (Shao et al., 2024)) as our optimization algorithm, which simplified value function of PPO (Schulman et al., 2017) with a normalized reward from a group. For the convention of notation, we use $\pi_{\theta}$ for $LM_{\text{SSA}}$ with parameter $\theta$. For completeness, we describe it here. It maximizes:
@@
Implementation We use a training set combining GSM8K and MATH (Cobbe et al., 2021; Hendrycks et al., 2021). To prepare the dataset, we use Qwen2.5-7B-Instruct model as $\text{LM}_{\text{ans}}$ to generate five answers for each questions (Qwen et al., 2025). Then we concatenate the question and each answer (ordered randomly) as input context of SSA LLM. We filter out instances where the combined answers exceed 4000 tokens, or where fewer than three valid answers are available. This results a dataset with a total size of 17.4k. We set the GRPO sample size to be 8, resulting in $8*17.4k\approx 140k$ during training.
@@ 3.3 Training SSA With Supervised Finetuning
Implementation We prompt Qwen 2.5 7B Instruct model for 5 candidate solutions per question in the GSM8K dataset. Then, we provide concatenated candidate solutions to GPT-4.1 Nano model along with the original question and ground-truth answer (The exact prompt is in Appendix B). GPT-4.1 Nano then provides a step-by-step reasoning process to identify and generate the best final answer from these candidates, yielding 7.47k training examples. Overall, this approach achieves a 96.24% match rate with the original ground-truth answers. [^1] When we inspected the remaining cases, we find many ground-truth labels are incorrect.
@@ 4.1 Experimental Settings
Datasets. For evaluation, we use an array of commonly used math reasoning datasets: the test split of GSM8K, MATH as the in-domain evaluation sets, and AIME 2024 (MAA, 2024), AMC 2023, and Olympiad (He et al., 2024) as the test sets. We use the extracted answers and grade them against the ground truth answers, using the library that has been used in prior work (Lightman et al., 2024).
Base Models for Candidate Generation ($\text{LM}_{\text{ans}}$). We use the Qwen-2.5-Instruct model, with sizes 7B, 14B, and 32B to generate $K$ of answers. We use a decoding temperature 0.5 to construct the training and test dataset. In the training phase, we use $k=5$ to train the SSA model. In the testing phase, we evaluated the performance with $k=\{5,10,15\}$.
Base Models for SSA ($\text{LM}_{\text{SSA}}$). We use the Qwen-2.5-base model with sizes 0.5B, 1.5B, and 3B. We chose the Qwen-2.5 model family due to its popularity and wide availability for PRM verifiers, making it possible to compare against existing PRM verifiers from the same model family.
Training Implementation Details. For the training library, we use torchtune due to its efficient VRAM management (torchtune maintainers & contributors, 2024). We use GRPO batch sample group size 8, batch size 1, temperature 1.0, AdamW optimizer, KL coefficient 0.01, and learning rate 1e-5. We trained all experiments one epoch. For hardware, we use 8*H100 80GB for training. For shorter context, it is also possible to train with 48GB VRAM.
@@ 4.2 Comparison Systems
Rule-based Baselines. We report pass@k, which reports the percentage of examples where any of $k$ number of solution is correct (Brown et al., 2025). Pass@k also serves as the oracle performance assuming we have an oracle verifier. We also report the majority vote (Wang et al., 2023), counting the most frequent answer among the answer set.
Outcome Reward Model (ORM). We use off the shelf Llama-3.1-8B ORM model trained with RLHF-Reward Modeling (Xiong et al., 2024) to re-rank multiple samples, selecting the best one scored by the reward. We will call it Llama-ORM (8B). It is trained with 273k data. For ORM evaluations, it adds an evaluation token at the end of the answer and the model will provide a score.
Process Reward Model (PRM). One prominent way to leverage multiple parallel sample is using process reward models. We use the Qwen-7B PRM (Qwen PRM) and Qwen-72B PRM model from Zhang et al. (2025c) to re-rank the candidate solutions.
@@
[…]
Universal Self Consistency (USC) (Chen et al., 2024a). USC is a prompting-based method takes a concatenation of the multiple parallel sample answers, and instructs a LM to generate the final answer. The exact prompt can be found in Appendix B. The task setting is equivalent to our SSA method, but the LM is not trained to optimize for the final answer.
SSA and Varitions. We train three types of the SSA models, one trained with the SFT objective only (SFT), one trained with SFT objective and then with the RL objective (SFT + RL), and one that is trained with RL objective only (RL). For SFT objective, we use the dataset described in Section 3.3 to finetune the model. We use learning rates 5e-5, 2e-5, and 1e-5 for 0.5B, 1.5B, and 3B correspondingly. For RL objective, we use GRPO method described in the Section 3.2. For SFT + RL objective, we use the trained model from the SFT objective, then we continue to train the RL objective on top of it. For each variant, we present models of three different sizes, trained from Qwen-0.5B, 1.5B, and 3B base models checkpoint respectively.
@@ 5 Main Results
Table 1: Results (accuracy %) with $k=5$ candidate answers generated by Qwen2.5-7B-Instruct as $\text{LLM}_{\text{ans}}$ model. Aggregation overheads (seconds/question) is measured end to end with AMC23 40 questions with k=5.
Method
Aggregation
Overhead (s)
Datasets Avg
@@
[1pt/1pt] USC w/ Qwen 3B 8.48 61.18 42.80 6.67 17.50 15.43 28.72
USC w/ Qwen 7B 5.89 5.89 61.20 6.67 47.50 28.78 43.83
LLama ORM (8B) 0.61 93.1 67.40 13.33 47.50 28.64 49.99
Qwen PRM (7B) 0.64 92.57 69.40 13.33 57.50 32.05 52.97
Qwen PRM (72B) 5.86 92.87 69.6 13.33 57.50 33.68 53.40
Shepherd PRM (7B) 0.65 90.75 64.40 13.33 35.00 27.00 46.10
[1pt/1pt] SSA RL (0.5B) 0.20 92.65 75.40 10.00 57.50 37.98 54.71
SSA RL (1.5B) 0.33 92.49 76.60 10.00 52.50 38.72 54.06
SSA RL (3B) 0.55 93.25 76.80 13.33 57.50 39.76 56.13